OpenStudy (anonymous):

I need help with a question and listening to Deadmau5 would help but it isn't pumping the brain power I thought it would. Solve the given equation. 6^3n+6=1,296

OpenStudy (anonymous):

i am going to make a guess that this is really \[6^{3n+6}=1,296\] am i right?

OpenStudy (anonymous):

here is why that is my guess: because \(1,296\) is an integer power of \(6\) that is \[6^4=1,296\]

OpenStudy (anonymous):

therefore you are being asked to sovle \[\huge 6^{\color{red}{3n+6}}=6^{\color{red}{4}}\]

OpenStudy (anonymous):

or "solve" whatever

OpenStudy (anonymous):

I mean the answer choices are this. http://prntscr.com/84jzo5 Sorry too lazy to type them out.

OpenStudy (anonymous):

You not solving using logs @satellite73

OpenStudy (anonymous):